OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!!

OpenStudy (abbycross167):

Please help me to simplify- x^2 - 3x - 4 -------------- x - 4

satellite73 (satellite73):

factor the numerator, then cancel the common factor of \(x-4\)

jimthompson5910 (jim_thompson5910):

to factor the numerator, find two numbers that... a) multiply to -4 (last term) AND b) add to -3 (middle coefficient) These two numbers will determine how the factorization goes

jimthompson5910 (jim_thompson5910):

For example x^2 + 5x + 6 factors to (x+3)(x+2) since 3*2 = 6 3+2 = 5

OpenStudy (abbycross167):

@jim_thompson5910 What do you mean find two numbers that... a) multiply to -4 (last term) AND b) add to -3 (middle coefficient)

jimthompson5910 (jim_thompson5910):

Do you see how x^2+5x+6 factors to (x+3)(x+2) ?

OpenStudy (abbycross167):

Yes sir

jimthompson5910 (jim_thompson5910):

notice how 3*2 = 6 and how 3+2 = 6

jimthompson5910 (jim_thompson5910):

so how would you factor something like x^2+10x+24 ?

OpenStudy (abbycross167):

Yes sir, I'm trying to figure it out now what two numbers multiple to -4

OpenStudy (abbycross167):

-2*2 = -4

jimthompson5910 (jim_thompson5910):

-2*2 = -4 is true but -2 plus 2 = 0 when we want the two numbers to add to -3

OpenStudy (abbycross167):

Yes sir I'm trying to figure that out, but I'm not very good with negatives

jimthompson5910 (jim_thompson5910):

what's another way to multiply to -4?

OpenStudy (abbycross167):

-4*1=-4 -4+1=-3 I didn't even think of this! The most simplest way!

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

the two numbers are -4 and +1 so x^2-3x-4 factors to (x-4)(x+1)

OpenStudy (abbycross167):

Ok then what?

jimthompson5910 (jim_thompson5910):

So we factor anything we can. Then we cancel out the common factors. In this case, a pair of (x-4) factors cancel out. \[\Large \frac{x^2 - 3x - 4}{x-4} = \frac{(x-4)(x+1)}{x-4}\] \[\Large \frac{x^2 - 3x - 4}{x-4} = \frac{\cancel{(x-4)}(x+1)}{\cancel{x-4}}\] \[\Large \frac{x^2 - 3x - 4}{x-4} = x+1\] where \[\Large x \ne 4\]

jimthompson5910 (jim_thompson5910):

So, \[\Large \frac{x^2 - 3x - 4}{x-4}\] simplifies to \[\Large x+1\] where \[\Large x \ne 4\]

OpenStudy (abbycross167):

Thank you so much!!! I appreciate your help!!